OpenStudy (anonymous):

In how many ways could you choose two different letters from the letters C, O, U, N, T? A. 60 ways B. 20 ways C. 120 ways D. 10 ways

OpenStudy (anonymous):

pretty sure it would be \[P_{5}^{2}\] so your equation would be \[\frac{ 5! }{ (5-2)! }\] meaning your answer would be 20

OpenStudy (anonymous):

Thank you! I have been trying to figure that out for the last 20mins. I knew it was P something

OpenStudy (anonymous):

there are 5 letters in total so out of 5 we choose anyone 4 the first place and now 4 2nd place there r 4 remaining so 4ways so total ways = 4*5=20

OpenStudy (anonymous):

i would say its \[\left(\begin{matrix}5 \\ 2\end{matrix}\right)\] = 10 ways since order doesnt matter right?

OpenStudy (anonymous):

we need 2 different letters

OpenStudy (anonymous):

Well, I did the 4 * 5 and I got 20, so Ima put it down and see.

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=20100520132011AAzJwcR this will help

OpenStudy (anonymous):

go for it @xDemonicWolfiex @jayz657 did u understand?

OpenStudy (anonymous):

Maybe it is 10. I dont want to miss this question. I know it, but I can't decide between 10 or 20

OpenStudy (anonymous):

i think you double counted since the order does not matter you get a double count ex: CO is the same as OC order doesnt matter so you dont count both

OpenStudy (anonymous):

@Mashy

OpenStudy (anonymous):

yea it should be 10 ways!

OpenStudy (anonymous):

yea was 90% sure since i am taking probability right now lol